OpenStudy (anonymous):

At a concession stand, three hotdogs and two hamburgers cost 8.50, two hotdogs and three hamburgers cost 9.00. What is the cost of one hamburger and one hotdog? So far I have: x=hotdog y-hamburger 3x+2y=8.50 2x+3y=9.00 For the life of me I cannot remember any of the steps past that. Any help would be awesome.

jimthompson5910 (jim_thompson5910):

Let's say we want to eliminate the y term. So we want the y coefficients to be equal but opposite in sign. This way, when we add the equations, they add to 0 and go away.

jimthompson5910 (jim_thompson5910):

To do this, multiply every term in the first equation by 3 to get 3x+2y=8.50 3(3x)+3(2y)=3(8.50) 9x+6y = 25.5

jimthompson5910 (jim_thompson5910):

Then multiply every term in the second equation by -2 2x+3y=9.00 -2*(2x)+(-2)*(3y)=-2*(9.00) -4x-6y = -18

jimthompson5910 (jim_thompson5910):

So we now have the equations 9x+6y = 25.5 -4x-6y = -18 Add them to get 9x+6y = 25.5 -4x-6y = -18 ------------- 5x+0y = 7.5 Which gives you 5x = 7.5

OpenStudy (anonymous):

Thank you so much. As soon as you started explaining it it call came flooding back! Much appreciated!

jimthompson5910 (jim_thompson5910):

that's great, glad it helped